Select the numeral preceding the expression that best answers
the question or completes the question
The lengths of the sides of a triangle may be
1) 4,9,13
2) 5,8,12
3) 5,5,13
4) 5,5,10

Found 2 solutions by LinnW, Edwin McCravy:
Answer by LinnW(1048)

Check each one.
1) 4,9,13
4^2+9^2 =? 13^2
16 + 81 not = 169 so this one does not work
2) 5,8,12
5^2 + 8^2 =? 12^2
25 + 64 not = 144 so this one does not work
3) 5,5,13
5^2 + 5^2 =? 13^2
25 + 25 not = 169 so this does not work
4) 5,5,10
5^2 + 5^2 =? 10^2
25 + 25 not = 100 so this does not work

Answer by Edwin McCravy(20081)

1) 4,9,13

Those three numbers CANNOT be the sides of a triangle because 
the two smaller numbers, 4 and 9, must have a sum which is
strictly greater than the third number.  4+9=13 and 13 is NOT
strictly greater than the third number 13. 


 2) 5,8,12

Those three numbers CAN be the sides of a triangle because 
the two smaller numbers, 5 and 8, have a sum which is
strictly greater than the third number.  5+8=13 and 13
is greater than the third number 12.

 3) 5,5,13

Those three numbers CANNOT be the sides of a triangle because 
the two smaller numbers, 5 and 5, must have a sum which is
strictly greater than the third number.  5+5=10 and 10 is NOT
greater than the third number 13. 

 4) 5,5,10

Those three numbers CANNOT be the sides of a triangle because 
the two smaller numbers, 5 and 5, must have a sum which is
strictly greater than the third number.  5+5=10 and 10 is NOT
strictly greater than the third number 10. 

Edwin